# YITP-26-13

### v1

@@ 3.4 Three-point functions and non-Gaussianity
$\langle\langle O(k_{1})O(k_{2})O(k_{3})\rangle\rangle_{\lambda_{2}}\propto\lambda_{2}k_{1}^{2}k_{2}^{2}\left(\frac{1}{E}+\frac{c_{3}k_{3}}{E^{2}}\right),$ (46)
@@ 5 Discussion and Summary
There are several promising directions for future research. First, the non-Gaussianity shapes derived here are relevant to the program of “Cosmological Collider Physics.” [75, 76, 77] The massive spin-1 or spin-0 modes in the aether sector could leave oscillatory signatures in the squeezed limit of the bispectrum, providing a unique probe of the bulk mass spectrum and symmetry-breaking pattern. Second, extending this analysis to theories with Hořava-Lifshitz scaling would be valuable. Since Einstein-Aether theory can be viewed as the low-energy limit of Hořava gravity, understanding the holographic renormalization group flow between these regimes would provide deeper insights into the emergence of spacetime and Lorentz invariance. Third, it would be interesting to revisit the effective field theory of vector-tensor theories [36] from the viewpoint of the distinction between scale and conformal invariance.
While the parameter space of the Einstein-Aether theory has been significantly constrained by the multi-messenger observation of GW170817 [63], binary pulsar data [78], and the latest LIGO-Virgo-KAGRA catalog [79], the theory remains a testing ground for Lorentz violation in gravity. Within the strong gravity regime, extensive studies have been conducted on black hole shadows and photon rings to confront the theory with Event Horizon Telescope observations [80, 81, 82], as well as on strong gravitational lensing effects [83, 84]. Dynamical features such as quasinormal modes and ringdown waveforms have also been rigorously computed to predict deviations from general relativity [85, 86]. Moreover, the theory’s consistency continues to be scrutinized through the first law of black hole thermodynamics [87] and positivity bounds derived from effective field theory principles [36, 88]. Our holographic analysis adds a new dimension to this active landscape by providing a dual field-theoretic perspective on the boost symmetry breaking.

### v2

@@ 3.4 Three-point functions and non-Gaussianity
$\langle\langle O_{1}(k_{1})O_{2}(k_{2})O_{3}(k_{3})\rangle\rangle_{\lambda_{2}}\propto\lambda_{2}k_{1}^{2}k_{2}^{2}\left(\frac{1}{E}+\frac{c_{3}k_{3}}{E^{2}}\right),$ (46)
@@ 5 Discussion and Summary
There are several promising directions for future research. First, the non-Gaussianity shapes derived here are relevant to the program of “Cosmological Collider Physics.” [75, 76, 77] The massive spin-1 or spin-0 modes in the aether sector could leave oscillatory signatures in the squeezed limit of the bispectrum, providing a unique probe of the bulk mass spectrum and symmetry-breaking pattern. Second, extending this analysis to theories with Hořava-Lifshitz scaling would be valuable. Since Einstein-Aether theory can be viewed as the low-energy limit of Hořava gravity, understanding the holographic renormalization group flow between these regimes would provide deeper insights into the emergence of spacetime and Lorentz invariance. Third, it would be interesting to revisit the effective field theory of vector-tensor theories [36, 78] from the viewpoint of the distinction between scale and conformal invariance.